A discriminatory farmer monopolist sells the product to college students and lecturers.


The demand function for students is Qsd=200-25p and the demand function for


lecturers is Qld=400-25p.Marginal cost is $1 per product. Calculate the amount of


product the monopolist sells for students and Lecturers. At what price level? [5


marks]


b. Suppose a firm is operating in a monopolistically competitive market structure.


i.Mention at least three characteristics of this market structure. [3 marks]


ii.What type of demand curve would the firm experience? Why? [2 marks]


iii.Draw the cost and revenue curves for a typical monopolistically competitive firm


in the short run for a given price level and explain how the firm chooses the level of


output that maximizes profit. [5 marks]

1
Expert's answer
2022-02-20T16:26:48-0500

Total quantity demanded"= Qsd+Qld"

"=200-25p+400-25p"

"Qd=600-50p"

"Total Revenue= Q\u00d7p= (600-50p)\u00d7p"

"=600p-50p^{2}"

Quantity supplied is maximum when Marginal revenue (M.R) equals (M.C).

"M.R= 600-100p"

Equating v"MR=MC"

"600-100p=1"

"100p=599"

"P=5.99"

"Qd=600-50\u00d7(5.99)"

"Qd= 300.5"

Characteristics of a monopolistic competition:

  • Relatively large number of sellers.
  • Defferentiated products.
  • Easy entry, easy exit.

Monopolistic competition has a kinked demand curve.